# Optimal Transportation by Orthogonal Coupling Dynamics

### v1

@@ 5. Representative Results for $L^{2}$-Monge-Kantorovich
In this section, we illustrate the accuracy and cost of the proposed OCD and highlight its capability in estimating the optimal map, along with Wasserstein distance, between two marginals in a few examples. Unless mentioned otherwise, by OCD we are referring to OCD-piecewise linear algorithm equipped with Runge-Kutta 4th order ODE solver. We refer the reader to Sec.7.4 in Supplementary Information for details on the choice of hyper-parameters such as $\epsilon$ and $\Delta t$, and comparison with benchmarks. All the performance studies are done on a single core and thread of Intel Core i7-8550U CPU with 1.80 GHz frequency.
@@ 5.1. Clustering and Prototypical Examples
Fist, we present and discuss three toy scenarios.
@@
    Symmetric Positive Semi-Definite Cone. As a by-product of OCD, we justified in Proposition 2.5 that the dynamics ensures the cross-correlation matrix remains in the cone of symmetric positive semi-definite matrices. Here we verify this finding by numerical simulation, where the initial condition is product measure of two centered Gaussians with covariance matrices $\Sigma_{\mu}$ and $\Sigma_{\nu}$. As shown in Fig. 4, OCD evolves the cross-correlation inside the cone and reaches the maximum on the curve connecting $\Sigma_{\mu}$ and $\Sigma_{\nu}$. Note that there exist interesting results on the geometry of the cross-correlation of Gaussians induced by the $L^{2}$-optimal transport and its connection to geometric average of $\Sigma_{\mu}$ and $\Sigma_{\nu}$, which we leave out of discussion for brevity, but refer the reader to [11].
@@
    Nonlinear Monge Map. To demonstrate relevance of OCD in recovering nonlinear multi-dimensional maps, we consider a synthetic setting where the target distribution is chosen such that the Monge map is known analytically. Consider the standard normal distribution over $\mathbb{R}^{2}$, denoted by $\mu$. Let the target distribution be the push-forward by softmax map $\nu=T\#\mu$, with $T=\nabla_{x}\log(\exp(x_{1})+\exp(x_{2}))$. Since $T$ is gradient of a convex, the Monge map between $\mu$ and $\nu$ is simply $T$. Depicted in Fig. 5, starting from independent $10^{6}$ samples of $\mu$ and $\nu$, particles pushed by OCD with $\epsilon=10^{-3}$ and $\Delta t=10^{-2}$ successfully recover the Monge map and accurately estimate 2-Wasserstein distance, as $t$ progresses.
@@ 5.2. Distribution Learning and Sampling
[…]
For testing, we generate $10^{6}$ normally distributed samples, i.e. $X^{\text{test}}\sim\mathcal{N}(0,I_{n\times n})$ and feed them to the NN to find $M_{X\rightarrow Y}(X^{\text{test}})$. As shown in Fig. 6, OCD is capable of recovering the target density with a reasonable accuracy. However, we note that the estimated densities by OCD are slightly smoothed out compared to the ground-truth. This error could be related to the current approach of estimating the conditional expectation, which resembles Kernel Density Estimation with known smoothing effects. In Sec. 7.4 of Supplementary Information, we provide a study of hyper-parameters and comparison with Adaptive Transport Maps (ATM) [8] for each test case. We point out that once the map is learned, computing the new samples is straight-forward, given that the network only performs simple matrix-vector multiplication on the samples of the standard normal distribution. Hence the proposed workflow can efficiently serve as a sampling approach for complex distributions (such as Swiss roll), with ramifications for sampling algorithms based on measure transport [60, 48].
@@
5.3. Classification in a Data Set
Since the Wassetstein distance provides a metric in the space of distributions, it can be leveraged to label data-sets in the context of machine-learning. For example, it allows for grouping similar data points and finding out their labels. To showcase the relevance of OCD in such applications, we compute the 2-Wasserstein distance in the Japanese Female Facial Expression (JAFFE) data-set [42]. As shown in Fig. 7, the proposed OCD (with $L^{2}$-cost) can find similarities between images. In particular, the distance between different images corresponding to the same model takes minimal values, on the diagonal elements. It is encouraging to see that given negligible memory requirement of OCD, we could carry out the computations on a local machine in a few seconds with fraction of memory requirement compared to EMD: for two $128\times 128$ pixel pictures, EMD consumes $6.5$ GB while OCD only needs $181$ MB of memory.
Refer to caption Refer to caption
Distribution of Wasserstein distance Example of found similar pictures
Figure 7. Distribution of the Wasserstein distance (left) and example of found similar pictures with minimum distance in each row (right) among $64\times 64$ pixels pictures in the Japanese Female Facial Expression (JAFFE) dataset using OCD with $\Delta t=0.1$ and $\epsilon=\epsilon_{\textrm{max}}(N_{\textrm{clusters}}/N_{p}>0.99)$.
@@ 5.4. Color Interpolation
As the final example, we consider another application of optimal transport in interpolation between color distribution of different images. Given an image with dimension $(N_{x},N_{y},3)$, the pixel colors can be considered as the random variable, leading to $N_{x}\times N_{y}$ samples for $X$ with $\dim(X)=3$, indicating red/green/blue channels. We take a low resolution picture of the ocean taken in day and sunset time for training, as depicted in Fig. 8 (see examples of POT library [29] for details).
@@
Given $N_{p}=10^{\prime}500$ training samples of these two pictures, we learn the map $M_{X\rightarrow Y}$ using paired samples found from OCD (with $L^{2}$-cost) and training NN similar to the one described in 5.2. We employ $\Delta t=0.1$ and $\epsilon=\epsilon_{\textrm{max}}(N_{\textrm{clusters}}/N_{p}>0.9)$. We repeat the same procedure for EMD to benchmark, and found that while EMD takes about $2.6$ GB of memory, OCD only consumes $115$ MB. Once the NN map is learned, we test the map by feeding the NN with the full resolution of day picture with $670\times 1000$ pixels, where we see a good agreement with the EMD solution. We further test the learned OCD map on two new testing picture with $548\times 1024$ and $1200\times 1920$ pixels. As shown in Fig. 9, the trained OCD map is capable of transforming the testing picture into sunset.
@@
Figure 9. Interpolating day and sunset pictures of ocean. The interpolation is done using the learned map $M_{X\rightarrow Y}(.)$ via $(1-\alpha)X+\alpha M_{X\rightarrow Y}(X)$. The bottom two rows show the transport of colors for two test picture.
@@ 6. Discussion
    The original problem of Monge-Kantorovich optimization, constrained on a set of couplings, is reduced to a dynamics whose evolution is controlled by the conditional expectation. The latter is solution of the regression problem, and thereby, the OCD provides a direct link between Bayesian regression and the optimal transportation.
@@
From numerical point of view, the model has favorable scaling behaviour. In the most basic form, it scales linearly with number of data points as well as dimension of distributional support, making OCD competitive in comparison to the state-of-the-art random map algorithms as well as direct linear programming approach. This powerful computational efficiency enables us to treat distributional learning problems over more complex and higher dimensional settings.
@@
Limitations. The current treatment of conditional expectation exhibits diffusive behaviour which can deteriorate marginal distributions with sharp features. More elaborated approaches, e.g. deployment of Gaussian Processes or Neural Nets as hypothesis class of the conditional expectation, may alleviate this issue, at least to some degree, though likely at the cost of increased computational complexity. In addition, moment constraints of the marginal distributions can be leveraged to enforce the marginal densities of OCD to remain close to the initial condition, and hence reducing the numerical dispersion. Another issue that requires further investigation is the optimal choice of $\epsilon$. Although we carried out extensive numerical tests, see Sec.  7.3 in Supplementary Information, we believe computing the conditional expectation efficiently is a separate task by itself. These trade-offs between accuracy and efficiency are interesting areas for future research on OCD. From a theoretical standpoint, our understanding of the underlying dynamics remains incomplete. While we provided some important theoretical results on OCD, questions such as the well-posedness of OCD and its convergence towards optimal solution are still unresolved and require further exploration.

### v2

@@ 6. Representative Results for $L^{2}$-Monge-Kantorovich
In this section, we illustrate the accuracy and cost of the proposed OCD and highlight its capability in estimating the optimal map, along with Wasserstein distance, between two marginals in a few examples. Unless mentioned otherwise, by OCD we are referring to OCD-piecewise linear algorithm equipped with Runge-Kutta 4th order ODE solver. We refer the reader to Supplementary Information for details on the choice of hyper-parameters such as $\epsilon$ and $\Delta t$, and comparison with benchmarks. All the performance studies are done on a single core and thread of Intel Core i7-8550U CPU with 1.80 GHz frequency.
@@ 6.1. Clustering and Prototypical Examples
First, we present and discuss three toy scenarios.
@@
    Symmetric Positive Semi-Definite Cone. As a by-product of OCD, we justified in Proposition 2.5 that the dynamics ensures the cross-correlation matrix remains in the cone of symmetric positive semi-definite matrices. Here we verify this finding by numerical simulation, where the initial condition is product measure of two centered Gaussians with covariance matrices $\Sigma_{\mu}$ and $\Sigma_{\nu}$. As shown in Fig. 4, OCD evolves the cross-correlation inside the cone and reaches the maximum on the curve connecting $\Sigma_{\mu}$ and $\Sigma_{\nu}$. Note that there exist interesting results on the geometry of the cross-correlation of Gaussians induced by the $L^{2}$-optimal transport and its connection to geometric average of $\Sigma_{\mu}$ and $\Sigma_{\nu}$, which we leave out of discussion for brevity, but refer the reader to [bhatia2019bures].
@@
    Nonlinear Monge Map. To demonstrate relevance of OCD in recovering nonlinear multi-dimensional maps, we consider a synthetic setting where the target distribution is chosen such that the Monge map is known analytically. Consider the standard normal distribution over $\mathbb{R}^{2}$, denoted by $\mu$. Let the target distribution be the push-forward by softmax map $\nu=T\#\mu$, with $T=\nabla_{x}\log(\exp(x_{1})+\exp(x_{2}))$. Since $T$ is gradient of a convex function, the Monge map between $\mu$ and $\nu$ is simply $T$. Depicted in Fig. 5, starting from independent $10^{6}$ samples of $\mu$ and $\nu$, particles pushed by OCD with $\epsilon=10^{-3}$ and $\Delta t=10^{-2}$ successfully recover the Monge map and accurately estimate 2-Wasserstein distance, as $t$ progresses. Note that the underestimation of the 2-Wasserstein distance is due to the fact that the marginals may not be preserved due to different sources of numerical errors, leading to, e.g. smaller variances of each marginal.
@@ 6.2. Distribution Learning and Sampling
We deploy the proposed OCD (with $L^{2}$-cost) to learn transport map between the normal source and a target density $\in$ $\{\textrm{Banana},\textrm{Funnel},\textrm{Swiss roll}\}$ [baptista2023representation]. We use $10^{4}$ samples of each marginal and carry out OCD to find the samples of the optimal joint density on $(\hat{X}^{*},\hat{Y}^{*})$. Afterwards, in order to learn the map via the paired data points, i.e. $M_{X\rightarrow Y}:\hat{X}^{*}\rightarrow\hat{Y}^{*}$, we deploy NN with $4$ layers, each with $100$ neurons, where Adam’s algorithm [diederik2014adam] with a learning rate of $0.002$ and $10^{4}$ iterations are utilized to find the NN weights.
For testing, we generate $10^{6}$ normally distributed samples, i.e. $X^{\text{test}}\sim\mathcal{N}(0,I_{n\times n})$ and feed them to the NN to find $M_{X\rightarrow Y}(X^{\text{test}})$. As shown in Fig. 6, OCD is capable of recovering the target density with a reasonable accuracy. However, we note that the estimated densities by OCD are slightly smoothed out compared to the ground-truth. This error could be related to the current approach of estimating the conditional expectation, which resembles Kernel Density Estimation with known smoothing effects. In Supplementary Information, we provide a study of hyper-parameters and comparison with Adaptive Transport Maps (ATM) [baptista2023representation] for each test case. We point out that once the map is learned, computing the new samples is straight-forward, given that the network only performs simple matrix-vector multiplication on the samples of the standard normal distribution. Hence the proposed workflow can efficiently serve as a sampling approach for complex distributions (such as Swiss roll), with ramifications for sampling algorithms based on measure transport [sun2024dynamical, papamakarios2021normalizing].
@@ 6.3. Color Interpolation
As the final example, we consider another application of optimal transport in interpolation between color distribution of different images. Given an image with dimension $(N_{x},N_{y},3)$, the pixel colors can be considered as the random variable, leading to $N_{x}\times N_{y}$ samples for $X$ with $\dim(X)=3$, indicating red/green/blue channels. We take a low resolution picture of the ocean taken in day and sunset time for training, as depicted in Fig. 7 (see examples of POT library [JMLR:v22:20-451] for details).
@@
Given $N_{p}=10^{\prime}500$ training samples of these two pictures, we learn the map $M_{X\rightarrow Y}$ using paired samples found from OCD (with $L^{2}$-cost) and training NN similar to the one described in section˜6.2. We employ $\Delta t=0.1$ and $\epsilon=\epsilon_{\textrm{max}}(N_{\textrm{clusters}}/N_{p}>0.9)$. We repeat the same procedure for EMD to benchmark, and found that while EMD takes about $2.6$ GB of memory, OCD only consumes $115$ MB. Once the NN map is learned, we test the map by feeding the NN with the full resolution of day picture with $670\times 1000$ pixels, where we see a good agreement with the EMD solution. We further test the learned OCD map on two new testing picture with $548\times 1024$ and $1200\times 1920$ pixels. As shown in Fig. 8, the trained OCD map is capable of transforming the testing picture into sunset.
6.4. Further numerical studies
In Supporting Information, we have provided more detailed numerical studies for the reader as a reference. This includes error analysis with respect to the OCD’s free parameters, and further comparisons against other benchmark methods including the invertible map [baptista2023representation] and Neural optimal transport [korotin2023neural]. Furthermore, we show solution of the OCD with the $L^{p}$-cost function as well as OT map between two 10-dimensional multivariate normal distributions.
@@ 6.4. Further numerical studies
Figure 8. Interpolating day and sunset pictures of ocean. The interpolation is done using the learned map $M_{X\rightarrow Y}(.)$ via $(1-\alpha)X+\alpha M_{X\rightarrow Y}(X)$. The bottom two rows show the transport of colors for two test picture. The histograms show the distribution of red, green, and blue colors for the corresponding image.
@@ 7.1. Concluding Remarks
    The original problem of Monge-Kantorovich optimization, constrained on a set of couplings, is reduced to a dynamics whose evolution is controlled by the conditional expectation. The latter is the solution of the regression problem, and thereby, the OCD provides a direct link between Bayesian regression and the optimal transportation.
@@
From numerical point of view, the model has favorable scaling behaviour. In the most basic form, it scales nearly linear with number of data points as well as dimension of distributional support, making OCD competitive in comparison to the state-of-the-art random map algorithms as well as direct linear programming approach. This powerful computational efficiency enables us to treat distributional learning problems over more complex and higher dimensional settings.
A crucial structural property of OCD, initialized by independent samples, is that the induced correlations remain symmetric and positive semidefinite. While this constitutes an interesting feature for randomized transport maps, it also implies that OCD, in its current deterministic formulation, does not in general recover the optimal coupling whenever the latter exhibits intrinsic asymmetries in its transport structure. In this sense, the symmetry constraint acts as an implicit bias, different in nature from the entropic bias induced by Schrödinger bridge or Sinkhorn-type methods (e.g. [leonard2013survey, cuturi2013sinkhorn, chen2021stochastic]). Whereas entropic algorithms favor couplings close to product measures, the OCD dynamics bias the evolution toward symmetric correlation structures.
The instability of suboptimal stationary solutions suggests a natural direction for relaxing this constraint: introducing controlled randomization into the dynamics, to explore the coupling space more flexibly. We anticipate that stochastic perturbations, e.g. in the form of annealing of the projected Langevin dynamics [conforti2023projected] may allow OCD-type flows to escape symmetric couplings and reach more general stationary couplings. A systematic theoretical characterization of the marginal distributions and structural conditions under which OCD recovers the optimal transport map remains an open research direction.
7.2. Limitations
@@
The current treatment of conditional expectation exhibits diffusive behaviour which can deteriorate marginal distributions with sharp features. More elaborated approaches, e.g. deployment of Gaussian Processes or Neural Nets as hypothesis class of the conditional expectation, may alleviate this issue, at least to some degree, though likely at the cost of increased computational complexity. In addition, moment constraints of the marginal distributions can be leveraged to enforce the marginal densities of OCD to remain close to the initial condition, and hence reducing the numerical dispersion. Another issue that requires further investigation is the optimal choice of $\epsilon$. Although we carried out extensive numerical tests, see Supplementary Information, we believe computing the conditional expectation efficiently is a separate task by itself. These trade-offs between accuracy and efficiency are interesting areas for future research on OCD. From a theoretical standpoint, our understanding of the underlying dynamics remains incomplete. While we provided some important theoretical results on OCD, questions such as the well-posedness of OCD and its convergence towards optimal solution are still unresolved and require further exploration.
7.3. Broader Context